EIGHTH INT’L PHONO MEETING/GP WORKSHOP, VIENNA, 2/11/96

Hungarian morphology doesn’t meet GP

Péter Rebrus^{1/3} <rebrus@nytud.hu>, Péter Siptár^{1/3} <siptar@nytud.hu>,

Péter Szigetvári^{2/3} <szigetva@nytud.hu> and Miklós Törkenczy^{1/2/3} <tork@nytud.hu>

[1] research institute for linguistics, hungarian academy of sciences (mta)

[2] english linguistics department, eötvös loránd university (elte)

[3] theoretical linguistics programme, elte/mta

CHART 1. Morphophonological alternations

nomin./gloss plural 2pl-poss. superess. inessive termin. -ness/-hood -like
type  stem and/or suffix does not change
zár ‘lock’  zárak  záratok  záron  zárban  zárig  zárság  zárszerű
i. LTL ól ‘pigsty’  ólak  ólatok  ólon  ólban  ólig  ólság  ólszerű
hegy ‘mountain’ hegy[�]k hegy[�]tek hegy[e]n hegyben hegyig hegység  hegyszerű
nyár ‘summer’ nyarak nyaratok nyáron nyárban nyárig nyárság nyárszerű

ii. SVS

tér ‘space’  terek  teretek  téren  térben  térig  térség  térszerű
iii. VA  mű ‘work’  művek művetek művön  műben  műig  műség  műszerű
bokor ‘bush’  bokrok bokrotok bokron  bokorban bokorig bokorság  bokorszerű
iv. VZA ökör ‘ox’  ökrök  ökrötök  ökrön  ökörben ökörig  ökörség  ökörszerű
selyem ‘silk’  selymek selymetek selymen  selyemben selyemig selyemség selyemszerű
anya ‘mother’  anyák anyátok anyán  anyában anyáig anyaság  anyaszerű

v. SFVL

remete ‘hermit’ remeték remetétek remetén  remetében remetéig remeteség remeteszerű
kár ‘damage’  károk  károtok  káron  kárban  kárig  kárság  kárszerű
kór ‘disease’  kórok  kórotok  kóron  kórban  kórig  kórság  kórszerű

vi. VH

tőr ‘dagger’  tőrök  tőrötök  tőrön  tőrben  tőrig  tőrség  tőrszerű
per ‘(law)suit’ per[e]k per[e]tek peren  perben  perig  perség  perszerű

stem and/or suffix changes

Note 1: Accent and double accent marks length, but a=[�], á=[a�], e=[�] (and [e] in the Dunántúl dialect),

é=[e�]; gy=[�], ly=[j], ny=[�], s=[�], sz=[s]. Note 2: [�]=[e] in the Budapest dialect. Note 3: The chart contains only nominal paradigms, and omits the accusative forms, which seem to be

regulated by some form of onset-to-onset government (the presence of presuffixal vowel is dependent on the sonority and/or place of articulation of the stem final consonant). All other nominal suffixes can be grouped with one of the types the columns represent.

Note 4: LTL = Lexically Triggered Lowering, SVS = Stem Vowel Shortening, VA = v-Augmentation, VZA =

Vowel–Zero Alternation, SFVL = Stem-Final Vowel Lengthening, VH = Vowel Harmony Note 5: There are 3 morphophonological alternation types which are available not only for closed stem-

classes: VH, SFVL, Predictable Lowering (i.e. lowering by non-lowering stems: verbal stems are never lowering, adjectival stems are almost exceptionlessly lowering, nominal stems vary with lowering stems being a closed class).

PROPOSITION 1. The system in Chart 1 reveals a hierarchy: there is an increasingly larger set of suffixes triggering the change; if n is triggered by a suffix, n + 1 is also triggered by it and if a suffix does not trigger n it does not trigger n − 1 either. PROPOSITION 2. Traditional GP analyses (Demirdache 1988, Dunn 1992, Polgárdi & Rebrus 1996, Rebrus 1996, Ritter 1995, Törkenczy 1992) have treated random sets of these alternations as phonological phenomena without giving any principled reason for the inclusion or exclusion of the others. PROPOSITION 3. Since i–iv in Chart 1 apply only to closed classes of stems they can be viewed as nonanalytical processes in the lexicon (Kaye 1995). This approach, however,

a. cannot explain the systematic property of suffixation mentioned in Proposition 1;

b. cannot explain the productivity of the suffixations which are unexplainable by phonotactic reasons: pun ‘Punic’ and punok ‘Punic, pl.’ vs. punk ‘punk’ bár ‘bar’ and bárod ‘your bar’ vs. bárd ‘hatchet, bard’

ANALYSIS 1: Synthetical suffixation. Alternation is regulated by proper government, SVS is seen as a subcase of VZA (Ritter 1995). Working hypothesis 1: presuffixal vowel is empty

1 bok@r⟨@⟩  nya@r⟨@⟩  ok  bokor, nyár
2 bok⟨@⟩r-@-k⟨@⟩  !nya⟨@⟩r-@-k⟨@⟩  melody problem for LTL bokrok, nyarak
3 bok⟨@⟩r-on⟨@⟩  *nya⟨@⟩r-on⟨@⟩  bokron, nyáron
4 *bok@r-⟨@⟩-tok⟨@⟩ *nya@r-⟨@⟩-tok⟨@⟩  bokrotok, nyaratok
5 *bok@r-⟨@⟩-k-@-t⟨@⟩ *nya@r-⟨@⟩-k-@-t⟨@⟩ plural accusative  bokrokat, nyarakat

Working hypothesis 2: presuffixal vowel is not empty

4′ bok⟨@⟩r-o-tok⟨@⟩ nya⟨@⟩r-a-tok⟨@⟩  bokrotok, nyaratok
5′ bok⟨@⟩r-o-k-a-t⟨@⟩ nya⟨@⟩r-a-k-a-t⟨@⟩  bokrokat, nyarakat

Problem: where does presuffixal vowel belong, to stem or to suffix?

a. to suffix, but then what is the difference between superessive and plural?

— either there are two plural suffixes, the choice between them lexically conditioned by stem and only one superessive

— or there is one plural suffix, but this is difficult to represent (this is not a VH process: tollak ‘pens, feathers’ vs. gólok ‘goals’ and dalok ‘songs’ vs. falak ‘walls’), it would need an element equivalent to [−low], which is unavailable, and the productive class would have to be marked lexically.

2′ bok⟨@⟩r-ok⟨@⟩  nya⟨@⟩r-ak⟨@⟩  bokrok, nyarak

b. to stem, but — domain-final full vowel has to disappear

— domain-final full vowel must not properly govern.

1′ *bok⟨@⟩ro  *nya⟨@⟩ra  bokok, nyár

A more sophisticated approach (László Kálmán p.c., Kornai 1992): floating A element:

1′′ bok@r⟨@⟩ nya@r⟨A⟩ bokor, nyár

but since all the elements of [�] (i.e., A, U and @) cannot be floating simultaneously, we need the following changes (also needed for the a∼á alternation in SFVL and SVS):

a language specific stipulation: .A → A.@

a universally plausible process: A.@ → AU.@ (nonlow back/non-A-headed vowels tend to be round) Before suffixes beginning with a nonempty vowel (e.g. superessive) the floating A deletes. Problems: needs ambient U (apparently needed anyway) and where does the [o] of bokrok come from (no default filling in in GP).

CONCLUSION 1. Since proper government in itself is inadequate to distinguish all stem types and melody is highly problematic, more than one level/degree of syntheticity need to be distinguished.

ANALYSIS 2: Analytical suffixation. Analyticity is revealed by morpheme-internally illicit C clusters. kádban ‘in the bathtub’: *[db] ⇒ [[kád]ban] or [[kád][ban]] kertben ‘in the garden’: *[rdb] ⇒ [[kert]ben] or [[kert][ben]]

It seems reasonable to distinguish harmonizing and nonharmonizing suffixes (e.g. -ban/-ben, -ság/-ség vs. -kor, -szerű) by positing the structures [[A]B] and [[A][B]], respectively. (Note that suffixes like -ig and -ért may be taken to be a harmonizing but contain transparent vowels.) Hence,

[[anyá]ban], [[anyá]ig], [[anya]ság] vs. [[anya][szerű]] SFVL cuts across this classification:

[[anyá]ban], [[anyá]ig] vs. anya∥ság, [[anya][szerű]]

CONCLUSION 2. Since VH groups anyában and anyaság together, while SFVL groups anyaság and anyaszerű together, a third type of analytic suffixation needs to be recognized.

references

Demirdache, Hamida. 1988. Transparent vowels. In: Harry van der Hulst and Norval Smith (edd.) Features, segmental

structure and harmony processes (Part II.). Dordrecht: Foris. 39–76. Dunn, Christian. 1992. Aspects du gouvernement harmonique. PhD dissertation. Université du Québec. Kaye, Jonathan. 1995. Derivations and interfaces. In: Jacques Durand and Francis Katamba (edd.) Frontiers of

phonology: Atoms, structures, derivations. Harlow: Longman. 289–332. Kornai, András. 1992. On Hungarian morphology. Budapest: Institutum Linguisticum Academiae Scientiarum Hun-

garicae. Polgárdi, Krisztina and Péter Rebrus. 1996. There is no labial harmony in Hungarian: A government phonology

approach. Working papers in the theory of grammar 3.3. Budapest: Research Institute for Linguistics, Hungarian Academy of Sciences.

Rebrus, Péter. 1996. Az analitikus–szintetikus toldalékolás a magyarban és a kormányzási fonológia [Analytical and synthetical suffixation in Hungarian and Government Phonology]. Néprajz és Nyelvtudomány [Ethnography and Linguistics] XXXVII : 141–156.

Ritter, Nancy Ann. 1995. The role of Universal Grammar in phonology: a Government Phonology approach to

Hungarian. PhD dissertation, New York University. Törkenczy, Miklós. 1992. Vowel–zero alternations in Hungarian: a government approach. In: István Kenesei and Csaba

Pléh (edd.) Approaches to Hungarian 4: The structure of Hungarian. Szeged: József Attila University (JATE). 157–